The average cost for a company to produce x units of a product is given by the function ​A(x)=(10x + 450)/x

Use A'(x) to estimate the change in average cost as production goes from 150 units to 151 units.

The change in average cost is approximately _____ dollars.
Answer #1

Average cost at 150 units = (10*150+450)/150 = $13 per unit

Average cost at 151 units = (10*151+450)/151 = $12.98 per unit

Change in average cost = $13-12.98 = $0.02 per unit (approximately)